rimare Corporation has provided the following data concerning last month’s manufacturing operations.

Purchases of raw materials $ 31,000
Indirect materials included in manufacturing overhead $ 4,820
Direct labor $ 58,500
Manufacturing overhead applied to work in process $ 88,900
Underapplied overhead $ 4,060
Inventories Beginning Ending
Raw materials $ 12,000 $ 18,500
Work in process $ 55,400 $ 65,300
Finished goods $ 34,900 $ 42,100

Required:

1. Prepare a schedule of cost of goods manufactured for the month.

2. Prepare a schedule of cost of goods sold for the month. Assume the underapplied or overapplied overhead is closed to Cost of Goods Sold.Primare Corporation has provided the following data concerning last months manufacturing operations. Purchases of raw materiComplete the question by entering your answers in the tabs below. Required 1 Required 2 Prepare a schedule of cost of goods s
Answer #1
Solution:
1st Primare Corporation
Schedule of cost of goods manufactured
Direct Materials:
a Beginning materials inventory 12,000
b Add: Purchases of raw materials 31,000
c=a+b Raw materials available for use 43,000
d Less: Ending raw materials inventory 18,500
e=c-d Raw materials used in production 24,500
f Less: Indirect materials included in manufacturing overhead cost incurred 4,820 19,680
g Direct Labor 58,500
h Manufacturing overhead applied to Work in Process 88,900
i=f+g+h Total manufacturing costs 167,080
j Add: Beginning work in process inventory 55,400
k=i+j 222,480
l Less: Ending work in process inventory 65,300
m=k-l Cost of goods manufactured 157,180
2nd.
Primare Corporation
Schedule of cost of goods sold
a Beginning finished goods inventory 34,900
b Add: Cost of goods manufactured 157,180
c=a+b cost of goods available for sale 192,080
d Less: Ending finished goods inventory 42,100
e=c-d Unadjusted cost of goods sold 149,980
f Add: Underapplied overhead 4,060
g=e+f Adjusted cost of goods sold 154,040
Notes: Underapplied overhead means applied overhead is lesser than actual overhead so it is added to unadjusted cost of goods sold to get accurate adjusted cost of goods sold . By adding this underapplied overhead we are allocating all the cost which are incurred for the goods sold.
Please feel free to ask if anything about above solution in comment section of the question.